OpenStudy (anonymous):

Does the ratio's form a proportion? 25/35 and 35/45???!

jimthompson5910 (jim_thompson5910):

They form a proportion if 25/35 = 35/45 is a true equation

jimthompson5910 (jim_thompson5910):

One way to check to see if 25/35 = 35/45 is a true equation is to cross multiply if you get both sides equal to the same number (after cross multiplication), then the original equation is true

OpenStudy (anonymous):

can you solve the variable? z/5 = 6/30 ??

jimthompson5910 (jim_thompson5910):

Example 1/2 = 3/6 1*6 = 2*3 ... cross multiply 6 = 6 ... this equation is true so 1/2 = 3/6 is true

jimthompson5910 (jim_thompson5910):

to solve z/5 = 6/30, we can cross multiply to get z*30 = 5*6 what's next?

OpenStudy (anonymous):

another solve the variable

OpenStudy (anonymous):

w/15 = 3/5

jimthompson5910 (jim_thompson5910):

same idea, hopefully the last one made sense?

OpenStudy (anonymous):

would the answer be 9?